The director of admissions at a large university says that 15% of high school juniors to whom she sends university literature eventually apply for admission. In a sample of 300 persons to whom materials were sent, 30 students applied for admission. In a two-tail test at the 0.05 level of significance, should we reject the director’s claim?


1
Expert's answer
2021-11-29T16:48:21-0500

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.15"

"H_1:p\not=0.15"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is"z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.1-0.15}{\sqrt{\dfrac{0.15(1-0.15)}{300}}}"

"\approx-2.4254"

Since it is observed that "|z| = 2.425 > 1.96=z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=2P(z<-2.4254)= 0.0152934," and since "p= 0.0152934<0.05," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is different than 0.15, at the "\alpha = 0.05"  significance level.